OpenStudy (anonymous):

Is (1,-4) a solution to the equation y=-2x a: yes b:no

jimthompson5910 (jim_thompson5910):

(1,-4) means that x = 1 and y = -4 This is because all points are of the form (x,y)

jimthompson5910 (jim_thompson5910):

Plug x = 1 and y = -4 into y=-2x y=-2x y=-2*1 ... replace x with 1 -4 = -2*1 ... replace y with -4 -4 = -2 ... multiply Since -4 = -2 is false, this means (1,-4) is NOT a solution to y=-2x So the point (1,-4) does NOT lie on the line y=-2x